In triangle WXY, WY=24m and angle Y=90degree. P lies on WX such that YP is perpendicular to WX, WP=18m and PX=14m. Q lies on YX such that QX=9.8m. Find
I) the length of YQ
ii) the area of triangle XPY

Answer by greenestamps(13200)



Triangles XPY and XYW are similar, because they are both right triangles that share angle X. So

PX/XY = XY/WX
(XY)^2 = PX*WX = 14*32 = 448
XY = sqrt(448) = 8*sqrt(7)

Then from the Pythagorean Theorem in triangle XPY, YP = 6*sqrt(7).

Answers...

(I) YQ = XY-XQ = 6*sqrt(7)-9.8
(II) the area of triangle XPY is (1/2)(PX)(PY) = (1/2)(14)(6*sqrt(7) = 42*sqrt(7)